Comparative financial statements for Weller Corporation, a merchandising company, for the year ending December 31 appear below. The company did not issue any new common stock during the year. A total of 600,000 shares of common stock were outstanding. The interest rate on the bonds, which were sold at their face value, was 10%. The income tax rate was 40% and the dividend per share of common stock was $0.40 this year. The market value of the company’s common stock at the end of the year was $30. All of the company’s sales are on account.

account.

Weller Corporation
Comparative Balance Sheet
(dollars in thousands)
This Year Last Year
Assets
Current assets:
Cash $ 1,210 $ 1,360
Accounts receivable, net 10,400 7,500
Inventory 12,400 11,000
Prepaid expenses 660 580
Total current assets 24,670 20,440
Property and equipment:
Land 9,300 9,300
Buildings and equipment, net 43,094 41,722
Total property and equipment 52,394 51,022
Total assets $ 77,064 $ 71,462
Liabilities and Stockholders' Equity
Current liabilities:
Accounts payable $ 18,900 $ 19,100
Accrued liabilities 1,080 770
Notes payable, short term 0 100
Total current liabilities 19,980 19,970
Long-term liabilities:
Bonds payable 8,800 8,800
Total liabilities 28,780 28,770
Stockholders' equity:
Common stock 2,000 2,000
Additional paid-in capital 4,000 4,000
Total paid-in capital 6,000 6,000
Retained earnings 42,284 36,692
Total stockholders' equity 48,284 42,692
Total liabilities and stockholders' equity $ 77,064 $ 71,462
Weller Corporation
Comparative Income Statement and Reconciliation
(dollars in thousands)
This Year Last Year
Sales $ 70,000 $ 65,000
Cost of goods sold 42,000 36,000
Gross margin 28,000 29,000
Selling and administrative expenses:
Selling expenses 10,600 10,900
Administrative expenses 6,800 6,800
Total selling and administrative expenses 17,400 17,700
Net operating income 10,600 11,300
Interest expense 880 880
Net income before taxes 9,720 10,420
Income taxes 3,888 4,168
Net income 5,832 6,252
Dividends to common stockholders 240 300
Net income added to retained earnings 5,592 5,952
Beginning retained earnings 36,692 30,740
Ending retained earnings $ 42,284 $ 36,692

Required:

Compute the following financial ratios for this year:

1. Times interest earned ratio.

2. Debt-to-equity ratio.

3. Equity multiplier.
Answer #1
1] Times interest earned ratio = EBIT/Interest expense = 10600/880 = 12.05
2] Debt to equity ratio = Total liabilities/Total stockholders' equity = 28780/48284 = 0.60
3] Equity multiplier = Total assets/Total stockholders' equity = 77064/48284 = 1.60